OpenStudy (anonymous):

Find the tr(A) such that (A^t- 5I)^-1= B where I is a 2 x 2 identity matrix where B is 0 1 -1 0

OpenStudy (anonymous):

sorry I asked this question but forgot that it was the inverse.

OpenStudy (freckles):

so if \[(A^t-5I)^{-1}=B \\\ (A^t-5I)^{-1}(A^t-5I)=B(A^t-5I) \\ I=B(A^t-5I)\] Let  so we have